An investment project has annual cash inflows of $4,200, $5,300, $6,100 and $7,400, and a discount rate of 14%, What is the discounted payback period for these cash flows if the initial cost is $7,000? What if the initial cost is $10,000? What if it is $13,000
Answer #1
I have already answered this question. You asked this question yesterday(28/2/12). Check it first.

Link is here
http://adf.ly/5szUs
Answer #2
When we use discounted payback, we need to find the value of all cash flows today. The value today of
the project cash flows for the first four years is:

Value today of Year 1 cash flow = $4,200/1.14 = $3,684.21
Value today of Year 2 cash flow = $5,300/1.142 = $4,078.18
Value today of Year 3 cash flow = $6,100/1.143 = $4,117.33
Value today of Year 4 cash flow = $7,400/1.144 = $4,381.39

To find the discounted payback, we use these values to find the payback period. The discounted first
year cash flow is $3,684.21, so the discounted payback for a $7,000 initial cost is:

Discounted payback = 1 + ($7,000 – 3,684.21)/$4,078.18 = 1.81 years

For an initial cost of $10,000, the discounted payback is:

Discounted payback = 2 + ($10,000 – 3,684.21 – 4,078.18)/$4,117.33 = 2.54 years

Notice the calculation of discounted payback. We know the payback period is between two and three
years, so we subtract the discounted values of the Year 1 and Year 2 cash flows from the initial cost.
This is the numerator, which is the discounted amount we still need to make to recover our initial
investment. We divide this amount by the discounted amount we will earn in Year 3 to get the fractional
portion of the discounted payback.

If the initial cost is $13,000, the discounted payback is:

Discounted payback = 3 + ($13,000 – 3,684.21 – 4,078.18 – 4,117.33) / $4,381.39 = 3.26 years
Answer #3
A = 4200 + 5300 + 6100 + 7400 = 23,000

1) 7000 - 23,000*0.14 = 3780

2) 10,000 - 23,000*0.14 = 6780

3)13,000 - 23,000*0.14 = 9780